OpenStudy (anonymous):

4. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x – 8y = 48 11x + 3y = -105.5

OpenStudy (precal):

no solution means they do not cross and are parallel infinite means they are the same line and overlap each other

OpenStudy (anonymous):

Multiply the first equation by 11, the second equation by -4 and add them up.

OpenStudy (precal):

yes that is one approach

OpenStudy (anonymous):

-11(4x-8y=48) 4(11x+3y=105.5) -44x+88y=-528 +44x+12y=422/100y=-106 Y=-1.06 4x-8(1.06)=48 4x+8.48=48 -8.48 -8.48 X=9.88<=4x=39.52

OpenStudy (precal):